OpenStudy (anonymous):

A garden contains two square flowerbeds .The total area of the flowebed is 680 square feet , and the second bed has 288 more square feet than the first . a) write a system of second-degree equation that models this situation. b) Graph the system found in part a and estimate the solution , c) Solve the system algebraically to find each flowerbed within the garden.

OpenStudy (anonymous):

Let me help you get started. For part a) We have two squares; thus, x^2 + y^2 = 680 and x^2 + 288 = y^2 Can you graph them?

OpenStudy (anonymous):

yes.Thanks